# When Should Selfish Miners Double-Spend?

### v3

@@ 1 Introduction
After Nakamoto’s seminal paper[1] that laid the foundations of blockchains aiming for large-scale consensus, there have been numerous works studying the related security guarantees and flaws [2]. The private attack, investigated initially in the Bitcoin whitepaper [1], aims to replace a transaction in the blockchain ledger with a conflicting transaction, allowing the attacker to double-spend its funds. Rosenfeld [3] showed rigorously that the private attack has exponentially decaying probability of success with time under the longest chain protocol. More specifically, as time passes, more public blocks are mined on top of a transaction that is being attacked, and it becomes less likely for a competing private chain to overtake the longest chain and undo the transaction in question.
@@
The celebrated selfish mining attack of Eyal and Sirer [4] exposed another vulnerability of the longest chain protocol by showing that it is not incentive-compatible. More specifically, [4] showed that a pool of miners, who constitute $\alpha$ fraction of the total hashrate in the network, can achieve a revenue ratio $\rho$, that is more than their fair share $\alpha$, i.e., the fraction of the hashrate they posses in the system. This is achieved by keeping the blocks they have mined private for a while according to the selfish mining algorithm and releasing them later to replace the blocks mined by honest miners, resulting in $\rho>\alpha$ in the long run. Moreover, the extra revenue ratio of selfish mining, i.e., $\rho-\alpha$, grows even further as the pool size increases, which attracts the rational miners to join the pool, eventually overtaking the chain completely.
@@
Later on, block withholding attacks and their effects on the revenue ratio have gained further attention and been studied extensively [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Sapirshtein et. al. [5] use Markov Decision Processes (MDP) to provide $\epsilon$-optimal policies for selfish mining attacks that result in higher revenue ratio than the algorithm of [4]. In a concurrent work, Nayak et. al. [6] provide an in-depth look at variations of selfish mining attacks, including the-now-famous stubborn mining, and analyze their revenue ratio using Monte Carlo simulation techniques. Further, Gervais et. al. [7] create an extended MDP problem where the model takes into account additional parameters such as the confirmation rule, network propagation dynamics, mining costs and eclipse attack. The authors of [7] further modify the model to consider the profitability of the double-spending attack via the MDP and investigate the impacts of the parameters on the optimal policies. Similar to the study of [7], [8] builds an MDP problem to investigate the profitability of the double-spending attack by slightly modifying the MDP model of [5] that replaces the rewards of override blocks by double-spend blocks defined therein. The authors in [8] further modify the MDP to investigate the strategies that optimally combine selfish mining attacks with double-spend attacks assuming a constant relation between the profits from the two attacks. The authors of [9] investigate a novel efficient approach to solve average reward ratio MDP of [5] using probabilistic termination methods. Other generalizations of the MDP model also exist that analyze selfish mining in different proof systems and protocols [10, 11, 12, 13, 14, 15]. Selfish mining strategies can further be combined with block withholding attacks on pools to increase revenue ratio [20, 21, 22]. Another line of research [16, 17] investigates the profitability of double-spending attacks using stochastic methods. Some of the recent works also focus on the profitability over time and use alternative metrics such as revenue per time or chain progress [23, 24, 25, 26].
@@
An important feature of selfish mining attack is the network influence of the pool, which determines the fraction of non-pool miners who favor the adversarial chain in case of fork-ties. This parameter, which is called $\gamma$, provides a threshold for the profitability of selfish mining, i.e., the minimum pool size needed for the selfish mining attack to be profitable. In other words, $\gamma$ determines the minimum $\alpha$ needed such that $\rho(\alpha,\gamma)>\alpha$. The authors of [4] argue that the current Bitcoin protocol, which picks the first chain seen by a miner in case of a tie, is vulnerable as the cost of increasing network influence in this case is minimal for the adversary. This in turn, resulted in search for new consensus policies that increase the threshold and the security [4, 27, 28, 29, 10]. Other studies provide methods to effectively calculate network influence $\gamma$ under various network models [30, 31]. We refer the interested reader to the survey in [32] for a further read on selfish mining and double-spending attacks and related defense mechanisms.
@@ 1.1 Our Contributions
Double-spending attacks in blockchain literature usually assume an adversary that does not focus on coinbase rewards and purely tries to undermine the consistency of the protocol. Such an analysis that ignores the adversarial revenue loss is usually not realistic. Hence, in this paper, for the first time in the literature, we provide a rigorous and explicit stochastic analysis of the combination of the two celebrated attacks: double-spending and selfish mining. Following the attack prescription we provide, gives the adversary repeated chances at double-spending without incurring revenue losses. To do so, we define a selfish mining attack strategy, called $L$-stubborn mining, that not only tracks the length difference between the pool’s private chain and the public chain but also its length. Tracking the length of the private chain allows us to determine whether the attack can result in double-spending or not in addition to the revenue steal from the honest miners. Essentially, $L$-stubborn mining, parametrized by the length of the private chain $L$, is a generalization of existing mining strategies such as honest mining ($L=1$), selfish mining ($L=2$), equal fork stubborn mining ($L=\infty$) where $L$ represents the stubbornness level of the attacker pool. Thus, an optimization of $L$ results in optimal revenue ratios $\rho_{L^{*}}$ among the no-trail stubborn strategies [6].
@@
More specifically, for each $\alpha$ and $\gamma$, we find the optimal $L^{*}$, i.e., the best strategy, under the condition that the adversary does not do any trailing behind, i.e., it accepts the honest branch whenever it falls behind. Similarly, we find largest $L$, that we call $\bar{L}$, such that $\rho>\alpha$, i.e., maximum stubbornness that still results in higher revenue ratio than honest mining strategy. Next, under the $k$-confirmation rule, we argue that if $\bar{L}>k$, then double-spending comes at no-cost. Hence, the results not only give a new understanding of selfish and stubborn mining but also provide parameter regimes of $\alpha$ and $\gamma$ where transactions are at risk from the perspective of double-spending. It is already known that when $\alpha\geq 0.5$ no transaction is safe. With the results we provide, for example for Bitcoin which follows $k$-confirmation rule with $k=6$, even if $\gamma=0$, every transaction is at risk of double-spend for $\alpha>0.409$. This, however, does not mean, that the transaction is going to be replaced with a conflicting one without fail, instead, it means, that trying the double-spending attack with no-trailing comes at no-cost to the adversary.
@@
We also provide a modified version of $L$-stubborn mining attack, which we call $S$-stealth mining, that is more similar to the classical private attack, in the sense that it does not expose itself by matching blocks and increases the success probability of double-spending attack at the cost of revenue. We analyze the properties of $S$-stealth mining as we do for the $L$-stubborn mining attack. For both $L$-stubborn mining and $S$-stealth mining attacks, we provide a new metric for revenue ratio calculation similar to the subversion gain defined in [10], where the extra revenue from the double-spending attack is taken into consideration in addition to the usual block rewards. Further, we also provide minimum transaction value for double-spending attacks that makes sure that the overall revenue ratio is larger than the honest mining strategy, when the usual revenue ratio is less than the honest mining strategy, similar to the subversion bounty defined in [10].
@@
  • Given pool size $\alpha$ and pool’s network influence $\gamma$, we find the optimal level of no-trail stubbornness $L$ and provide the associated revenue ratio.
@@
Different from [7, 8], we provide a simple formula to calculate revenue ratios, double-spend risks and optimal attacks. As we do not deal with MDPs, the numerical results in Section 6 are extremely easy-to-analyze and provide a great insight at first sight. Moreover, the strategy we provide is easy-to-follow compared to complicated prescriptions of MDPs. Our results can directly be used to rigorously analyze optimal mining revenue per time, a recent focus of research topic [23, 24, 25, 26].
@@
Our paper is structured as follows. In Section 2, we provide the essential probabilistic tools in a concise manner that are needed to understand the analysis provided in this paper. In Section 3, we define the system model, provide the attackers’ mining strategies and relevant definitions. In Section 4, we rigorously analyze the revenue ratio of the $L$-stubborn attack as well as the double-spending that comes with it. We do the same for $S$-stealth mining strategy in Section 5. In Section 6, we provide the relevant numerical results. Finally, in Section 7, we discuss the severity and potential implications of the attack as well as possible extensions.
2 Preliminaries
2.1 Catalan Numbers and Dyck Words
$C_{n}=\frac{1}{n+1}{2n\choose n}$. (1)
Catalan numbers also satisfy the following recursion
$C_{n}=\frac{2(2n-1)}{n+1}C_{n-1}$ (2)
with $C_{0}=1$. The generating function of the Catalan numbers is
$C(x)=\sum_{n\geq 0}C_{n}x^{n}=\frac{2}{1+\sqrt{1-4x}}.$ (3)
Every Pre-Dyck word starts with $X$ and every Dyck word ends with $Y$, i.e., $P[n,n]=P[n,n-1]=C_{n}$. A Pre-Dyck word can end with $Y$ if $m>0$ and with $X$ if $n>m$. Thus, Pre-Dyck words satisfy the recursion
$P[n,m]=P[n,m-1]+P[n-1,m].$ (4)
Clearly $P[n,m]=0$ if $m>n$ and it is easy to verify that $P[n,0]=1$ and $P[n,1]=n$. Thus, (4) gives a straightforward formula for finding $P[n,m]$ in $\mathcal{O}(nm)$. Every Dyck word with length $2n$ has a suffix that starts with 11 $X$ and $r$ $Y$’s, which implies
$C_{n}=\sum_{m=0}^{n-1}P[n-1,m].$ (5)
2.2 Bailey’s Number and Bertrand’s Ballot Problem
$P[n,m]=\frac{n-m+1}{n+m+1}{n+m+1\choose n+1}$. (6)
Note that, if a broad range of $P[n,m]$ is needed, it is better to start from $P[1,0]=1$ and fill the matrix $P$ via (4). The elements of the matrix $P$ constitute Catalan’s triangle.
The elements of $P$ matrix appear in [34] where Bailey investigates the sequences that contain $n$ 11’s and $m$ 1-1’s such that the running sum is always non-negative. Bailey proved that the number of such sequences is equal to $P[n,m]$ provided in (6). Further, the diagoanal and antidiagonal (including all the below diagonal and antidiagonal) elements of matrix $P$ overlap with the columns of $A^{(1,2)}$ defined in [41] as Catalan-like numbers, which is not surprising since the recursion equations are similar.
2.3 Wald’s Identity and Doob’s Optional Stopping Time
@@ 3 System Model
We assume a blockchain running under the Nakamoto consensus, i.e., PoW with longest chain protocol. The mining model we consider is the same as the model used in the seminal work on selfish mining by Eyal and Sirer[4] and the subsequent works on the topic [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. More specifically, we assume that there are three groups of miners. The first group can be considered as a single-entity which constitutes $\alpha=1-\beta$ fraction of the mining power in the system and acts adversarially, i.e., does not follow the longest-chain protocol and can decide to mine privately without sharing the blocks it has mined. In this paper, we call this entity as the adversary. The second and third groups together comprise the honest miners who follow the longest-chain protocol, i.e., mine on top of the longest-chain they have seen and release the newly mined blocks immediately. The former of the two constitutes $\gamma\beta$ fraction of the mining power is influenced by the adversarial group and accepts the adversarial chain in case of a tie, whereas the latter constitutes $(1-\gamma)\beta$ fraction of the mining power and accepts the honest chain in case of a tie. The reason for such a partition and in which real-world scenarios this can arise is explained in [4].
@@
We only consider the blocks that contain a valid PoW (Proof-of-Work) and assume that the inter-arrival times between blocks follow an exponential distribution. When a block is mined by an honest miner, it is subject to a propagation delay. During this propagation delay, if the adversary releases a competing block at the same height, called matching, $\gamma$ fraction of the honest miners receive the adversarial block first and mine on top of it. On the other hand, we ignore the newly mined blocks during the propagation delay. As a result, each mined block belongs to one of the three groups with a probability proportional to their mining power.
@@
Moreover, the blockchain protocol updates the puzzle difficulty threshold so that the average inter-arrival time between blocks is constant. Each block rewards its miner a predetermined amount of coins, i.e., coinbase, which we assume is constant for the foreseeable future. If each miner acts honestly, the coinbase reward they get is proportional to their mining power in the network in the limit. However, adversarial strategies which deviate from the longest-chain protocol result in rewards disproportional to the mining powers. We refer to the adversarial strategies that aim to increase their fraction of the coinbase rewards as selfish mining strategies. Although these selfish mining strategies initially decrease the rewards, they can be seen as an attack on the block interval adjustment [42], which in turn increases the rewards in the long run.
A transaction $tx$ in this system is considered to be confirmed according to a $k$-deep confirmation rule if a block that contains $tx$ has k1k-1 blocks mined on top of it and is part of the longest chain available to the honest miners. Double-spending refers to the event that $tx$ is confirmed according to the $k$-deep confirmation rule and a conflicting transaction $tx^{\prime}$ replaces $tx$ in the longest chain and becomes $k$-deep.

### v4

@@ 1 Introduction
After Nakamoto’s seminal paper[1] that laid the foundations of blockchains aiming for large-scale consensus, there have been numerous works studying the related security guarantees and flaws [2]. A rule-of-thumb in Nakamoto blockchains is to treat a block that is $k$-block deep in the longest chain as a point of consensus, i.e., consider it as confirmed (e.g., $k=6$ for Bitcoin). The private attack, investigated initially in the Bitcoin whitepaper [1], aims to replace a confirmed block (hence, the transactions within) with a conflicting one, allowing the attacker to double-spend its funds. Rosenfeld [3] showed rigorously that the private attack has exponentially decaying probability of success with time under the longest chain protocol. More specifically, as time passes, more public blocks are mined on top of a block that is being attacked, and it becomes less likely for a competing private chain to overtake the longest chain and undo the block in question. In that sense, a block that is sufficiently deep in the longest chain is a point of consensus and double-spending is a poor operating procedure for a fixed $k$ of the $k$-confirmation rule which is a heuristic approach to reach consensus.
@@
The celebrated selfish mining attack of Eyal and Sirer [4] exposed another vulnerability of the longest chain protocol by showing that it is not incentive-compatible. More specifically, [4] showed that an adversarial miner, who possesses $\alpha$ fraction of the total hashrate in the network, can achieve a revenue ratio $\rho$, that is more than its fair share $\alpha$. This is achieved by keeping the blocks it has mined private for a while according to the selfish mining algorithm and releasing them later to replace the blocks mined by honest miners, resulting in $\rho>\alpha$ in the long run. Moreover, the extra revenue ratio of selfish mining, i.e., $\rho-\alpha$, grows even further as the adversarial hash fraction in the system increases, which attracts the rational miners to join the adversary, eventually overtaking the chain completely.
@@
Later on, block withholding attacks and their effects on the revenue ratio have gained further attention and been studied extensively [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Sapirshtein et. al. [5] use Markov Decision Processes (MDP) to provide $\epsilon$-optimal policies for selfish mining attacks that result in higher revenue ratio than the algorithm of [4]. In a concurrent work, Nayak et. al. [6] provide an in-depth look at variations of selfish mining attacks, including the-now-famous stubborn mining as well as its no-trail variations, and analyze their revenue ratio using Monte Carlo simulation techniques. Further, Gervais et. al. [7] create an extended MDP problem where the model takes into account additional parameters such as the confirmation rule, network propagation dynamics, mining costs and eclipse attack. The authors of [7] further modify the model to consider the profitability of the double-spending attack via the MDP and investigate the impacts of the parameters on the optimal policies. Similar to the study of [7], [8] builds an MDP problem to investigate the profitability of the double-spending attack by slightly modifying the MDP model of [5] that replaces the rewards of override blocks by double-spend blocks defined therein. The authors in [8] further modify the MDP to investigate the strategies that optimally combine selfish mining attacks with double-spend attacks assuming a constant relation between the profits from the two attacks. The authors of [9] investigate a novel efficient approach to solve average reward ratio MDP of [5] using probabilistic termination methods. Other generalizations of the MDP model also exist that analyze selfish mining in different proof systems and protocols [10, 11, 12, 13, 14, 15]. Selfish mining strategies can further be combined with block withholding attacks on pools to increase revenue ratio [20, 21, 22]. Another line of research [16, 17] investigates the profitability of double-spending attacks using stochastic methods. Some of the recent works also focus on the profitability over time and use alternative metrics such as revenue per time or chain progress [23, 24, 25, 26].
@@
An important feature of selfish mining attack is the network influence of the adversary, which determines the fraction of honest miners who favor the adversarial chain in case of fork-ties. This parameter, which is called $\gamma$, provides a threshold for the profitability of selfish mining, i.e., the minimum adversarial hash fraction needed for the selfish mining attack to be profitable. In other words, $\gamma$ determines the minimum $\alpha$ needed such that $\rho(\alpha,\gamma)>\alpha$. The authors of [4] argue that the current Bitcoin protocol, which picks the first chain seen by a miner in case of a tie, is vulnerable as the cost of increasing network influence is minimal for the adversary. This in turn, resulted in search for new consensus policies that increase the threshold and the security [4, 27, 28, 29, 10]. Other studies provide methods to effectively calculate network influence $\gamma$ under various network models [30, 31]. We refer the interested reader to the survey in [32] for a further read on selfish mining and double-spending attacks and related defense mechanisms.
@@ 1.1 Our Contributions
Double-spending attacks in blockchain literature usually assume an adversary that does not focus on coinbase rewards and purely tries to undermine the consistency of the protocol. Such an analysis that ignores the adversarial revenue loss is usually not realistic. Hence, in this paper, for the first time in the literature, we provide a rigorous and explicit stochastic analysis of the combination of the two celebrated attacks: double-spending and selfish mining. Following the attack prescription we provide, gives the adversary repeated chances at double-spending without incurring revenue losses. To do so, we define a selfish mining attack strategy, called $L$-stubborn mining, that not only tracks the length difference between the adversary’s private chain and the public chain but also its length. Tracking the length of the private chain allows us to determine whether the attack can result in double-spending or not in addition to the revenue stolen from the honest miners.
Essentially, in $L$-stubborn mining, the adversary keeps mining on its chain privately until it reaches length $L$ as long as it does not fall behind of the public honest chain. From this point onwards, the adversary does not risk falling behind the public honest chain, hence releases its private chain if and when the honest chain is one block behind. Note that, parametrized by $L$, $L$-stubborn mining is a generalization of existing mining strategies in the literature, such as honest mining ($L=1$), selfish mining ($L=2$), equal fork stubborn mining ($L=\infty$) where $L$ represents the stubbornness level of the adversary. Thus, an optimization of $1\leq L\leq\infty$ results in optimal revenue ratios $\rho_{L^{*}}$ among the no-trail stubborn strategies.
@@
More specifically, for each $\alpha$ and $\gamma$, we find the optimal $L^{*}$, i.e., the best strategy, under the condition that the adversary does not do any trailing behind, i.e., it accepts the honest branch whenever it falls behind. Similarly, we find largest $L$, that we call $\bar{L}$, such that $\rho>\alpha$, i.e., maximum stubbornness that still results in higher revenue ratio than honest mining strategy. Next, using the $k$-confirmation rule as a heuristic to reach consensus as is the convention, we argue that if $\bar{L}>k$, then double-spending comes at no-cost. Hence, the results not only give a new understanding of selfish and stubborn mining but also provide parameter regimes of $\alpha$ and $\gamma$ where transactions are at risk from the perspective of double-spending. It is already known that when $\alpha>0.5$ no transaction is safe. With the results we provide, for example for Bitcoin which follows $k$-confirmation rule with $k=6$, even if $\gamma=0$, every transaction is at risk of double-spend for $\alpha>0.409$. This, however, does not mean, that the transaction is going to be replaced with a conflicting one without fail, instead, it means, that trying the double-spending attack with no-trailing comes at no-cost to the adversary.
@@
We also provide a modified version of $L$-stubborn mining attack, which we call $S$-stealth mining, that is more similar to the classical private attack, in the sense that it does not expose itself by matching blocks and increases the success probability of double-spending attack at the cost of revenue. We analyze the properties of $S$-stealth mining as we do for the $L$-stubborn mining attack.
@@
  • Given the adversarial hash fraction $\alpha$ and adversarial network influence $\gamma$, we find the optimal level of no-trail stubbornness $L$ and provide the associated revenue ratio.
@@
Different from [7, 8], we provide a simple formula to calculate revenue ratios, double-spend risks and optimal attacks. As we do not deal with MDPs, the numerical results in Section 6 are extremely easy-to-analyze and provide a great insight at first sight. Moreover, the strategy we provide is easy-to-follow compared to complicated prescriptions of MDPs. We numerically compare $\rho_{L^{*}}$ with the revenue ratio of the $\epsilon$-optimal selfish mining attack to argue that our strategy performs close to the optimal solution of the MDP of [5]. Hence, our results can be a substitute to rigorously analyze optimal mining revenue per time, a recent focus of research topic, where MDP solutions cannot be used directly [23, 24, 25, 26].
@@
Our paper is structured as follows. In Section 2, we define the system model, provide the attackers’ mining strategies and relevant definitions. In Section 4, we rigorously analyze the revenue ratio of the $L$-stubborn attack as well as the double-spending that comes with it. We do the same for $S$-stealth mining strategy in Section 5. In Section 6, we provide the relevant numerical results. Finally, in Section 7, we discuss the severity and potential implications of the attack as well as possible extensions.
@@ 2 System Model
We assume a blockchain running under the Nakamoto consensus, i.e., PoW with longest chain protocol. The mining model we consider is the same as the model used in the seminal work on selfish mining by Eyal and Sirer[4] and the subsequent works on the topic [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. More specifically, we assume that there are two groups of miners. The first group can be considered as a single-entity which constitutes $\alpha=1-\beta<0.5$ fraction of the mining power in the system and acts adversarially, i.e., does not follow the longest-chain protocol and can decide to mine privately without sharing the blocks it has mined. In this paper, we call this entity as the adversary. The second group comprises the honest miners who follow the longest-chain protocol, i.e., mine on top of the longest-chain they have seen and release the newly mined blocks immediately. In case of a tie, the honest miners prefer the chain they first became aware of.
@@
In this paper, we only consider the blocks that contain a valid PoW (Proof-of-Work) and call the blocks mined by an honest miner as an honest block whereas a block with a valid PoW mined by the adversary is called an adversarial block. Here, we note that the inter-arrival times between blocks follow an exponential distribution due to the properties of PoW mining. When a block is mined by an honest miner, it is subject to a small propagation delay. During this propagation delay, if the adversary has a competing block at the same height that is kept private, it can rush to release the block. In this context, releasing a block means that the adversary shares the block with the rest of the network, i.e., the honest miners. This action is called matching, and subsequently, $\gamma$ fraction of the honest miners receive the adversarial block first and mine on top of it, whereas the rest of the honest miners receive the honest block first and mine on top of it. Hence, $\gamma$ is called the adversarial network influence or rushing factor. We assume that the adversary is aware of any block in the network as soon as that block is mined. Any block that is known/observed by honest miners is called public block and the public chain refers to the longest chain that exclusively consists of public blocks. However, during a fork where the adversary matches an honest block on the tip of the longest chain by releasing its private block at the same height, for the sake of distinguishing the two forks that are public and of equal length, we call the branch released by the adversary as the adversarial branch even though it is public as well.
@@
The blockchain protocol updates the puzzle difficulty threshold so that the average inter-arrival time between blocks is constant. Each block rewards its miner a predetermined amount of coins, i.e., coinbase, which we assume is constant for the foreseeable future. If each miner acts honestly, the coinbase reward they get is proportional to their mining power in the network in the limit. However, adversarial strategies which deviate from the longest-chain protocol result in rewards disproportional to the mining powers. We refer to the adversarial strategies that aim to increase their fraction of the coinbase rewards as selfish mining strategies. Although these selfish mining strategies initially decrease the rewards, they can be seen as an attack on the block interval adjustment [40], which in turn increases the rewards after the difficulty is adjusted.
In this work, we use the heuristic of $k$-confirmation rule for reaching consensus. A block in this system is considered to be confirmed according to a $k$-deep confirmation rule if it is part of a longest chain available to the honest miners. Instead of defining and using conflicting $\mathsf{tx}$s to refer to double-spending events, in this paper, we simply use the notion of multiple blocks at the same height being confirmed (at the same time or at two different points in time) as double-spending events. The heuristic of $k$-confirmation rule for reaching consensus is the convention in the literature and part of our system model where we pick a fixed $k$ to analyze the double-spending event. This is because honest miners are assumed to not afford waiting too long to confirm a block. However, it is important to mention that if latency to confirm a block is not of importance to honest miners, they could avoid the poor operating procedure of double-spending by waiting sufficiently long before confirming a block.